# Shot-to-shot acquisition ultrafast electron diffraction

### v1

@@ III.2 Noise level with pump
In the presence of the pump, other noise sources from the thermal and mechanical fluctuation of the thin film should be considered. To verify how it affects real measurement, we focus our treatment on a pump-probe experiment before the temporal overlap between the pump and the probe, i.e. before $t_{0}$. The reversible signal after the photo-excitation won’t interfere with our treatment, but only the long-lived thermal and mechanical fluctuations.
We take a 50 nm thick sample of natural graphite as a studied specimen whose diffraction patterns are shown in Fig. 6c). We photoexcite the sample with an energy density of 0.8 mJ/cm2 at a wavelength around 800 nm. Similar experiments have been reported with ultrafast electron diffraction Carbone et al. (2008); Chatelain et al. (2014); Stern et al. (2018), but in this section, we analyze the signal before the arrival time of the photoexcitation pulse, which should correspond to the diffraction patterns without photoexcitation in the case of a reversible photoexcitation.
We analyze a set of $I_{\text{on}}^{(i)}$ and $I_{\text{off}}^{(j)}$ images corresponding to a ten-second exposure of graphite’s diffraction patterns when it is photoexcited for a time window $i$, and when it is at rest for a time window $j$, respectively. For each diffraction image, we used a mask to set the unscattered electron beam to zero such that we compare only the fluctuation of the scattered electron. Despite a protection coating on the detector, the pump beam scattered in the chamber creates abnormal high-intensity pixels that are only visible on the $I_{\text{on}}$ images. To remove them, we compare each image $I_{\text{on}}^{(i)}$ to $I_{\text{off}}^{(i)}$. The pixels on $I_{\text{on}}^{(i)}$ respecting the conditions $I_{\text{on}}^{(i)}/I_{\text{off}}^{(i)}>n$ and $I_{\text{on}}^{(i)}/I_{\text{off}}^{(i)}<1/n$ are set to their $I_{\text{off}}^{(i)}$ value. For a cutting parameter $n=2$, around $1.8\%$ of the pixels are neutralized.
Fig. 3a) shows the correlation between $I_{\text{on}}^{(i)}$ and $I_{\text{off}}^{(i)}$ taken shot-to-shot, and between two diffraction patterns $I_{\text{on}}^{(i)}$ and $I_{\text{on}}^{(j)}$ taken few seconds apart, i.e. $i\neq j$. The pulse-to-pulse images are strongly correlated compared to the two single images, especially at low electron count pixel. We can characterize the correlation between the two images with the Pearson coefficient:
@@
where the sum is over all the pixel $x$ and $\langle I\rangle$ corresponds to the mean over the pixels. Fig 3b) shows the Pearson coefficient for a set of integrated images taken at different times. The correlation between $I_{\text{on}}^{(i)}$ and $I_{\text{off}}^{(i)}$ images taken shot-to-shot stays close to 1 while the one between $I_{\text{on}}^{(0)}$ and $I_{\text{on}}^{(j)}$ decreases and fluctuates with the time span between exposure. The average correlation between $\langle I_{\text{on}}(i),I_{\text{off}}(i)\rangle_{i}$ is $6.8\cdot 10^{-4}$ close to 1 with a standard deviation of $1.6\cdot 10^{-5}$. It suggests that the shot-to-shot acquisition mode is resilient to thermal and mechanical fluctuation, and these fluctuations stay stable.
@@
Fig 3c) shows the pixel distribution of the normalization $I_{\text{on}}/I_{\text{off}}$ as a function of exposure time. We observe that it converges towards a Gaussian distribution with decreasing variance reaching $0.01$ with a mean of $0.99$. It confirms the robustness of the correlation between the diffraction pattern of the unpumped and pumped graphite.
Figure 4: Transient signal due to plasma generation after photoexcitation on a copper grid. Zoom in the unfocused electron beam that went through the copper grid after the photoexcitation photon pulse in a) the presence of the pump $I_{\text{on}}$, b) the absence of the pump $I_{\text{off}}$, and c) the normalization between each other $I_{\text{on}}/I_{\text{off}}$. d) Comparison between the transient signals measured for a region of interest in the center of the electron beam $I_{\text{on}}(t)$, $I_{\text{off}}(t)$, and $I_{\text{on}}(t)/I_{\text{off}}(t)$. e) Comparison of the normalized shot-to-shot transient signal of the plasma generation on the copper grid for different electron fluxes fitted with equation 4.
III.3 Time resolution
One of the most important features of a UED setup is the time duration of the electron pulse at the sample position which defines the temporal resolution of the experimentHassan et al. (2017). The space charge effect broadening the electron pulse in time and space is the limiting factor of the time resolution. To limit the space charge effect without a temporal compression scheme such as RF cavity or THz fieldChatelain et al. (2012); Ehberger et al. (2019), we place the electron source as close to the sample position as mechanically possible - around $15$ cm.
We characterize the pulse duration to some extent with the convolution of step-function plasma generation $\Psi(t)$ on a copper grid and the electron pulse Gaussian distribution $g(t)$:
$f(t)=\Psi(t)\star g(t)=A\cdot\text{Erfc}\left(\frac{t-t_{0}}{\sqrt{2}\sigma}\right),$ (4)
with $t_{0}$ the time when the pump and the probe overlap on the sample, and $A$ the signal amplitude. The parameter $\sigma$, corresponding to the variance of the Gaussian electron distribution, defines the temporal resolution of the experiment at a given electron flux.
We use a 1000 mesh copper grid and photoexcite at a wavelength of 800 nm with a fluence of 2 mJ/cm2. Fig. 4a) and b) show the unfocused electron beam that went through the copper grid in the absence, $I_{\text{on}}(t)$, and the presence, $I_{\text{off}}(t)$, of the pump respectively at time $t$ after the photoexcitation. Thanks to the shot-to-shot acquisition detection, the normalization between Fig. 4a) and b), $I_{\text{on}}/I_{\text{off}}$, shown in Fig. 4c), clearly reveals the pattern of the plasma after the photoexcitation.
Fig. 4d) shows the total count in a region of interest defined as a $6\times 6$ pixels highlighted by the square in Fig. 4c). Each delay corresponds to an exposure time $20\cdot 10^{3}$ pulses for both $I_{\text{on}}(t)$, and $I_{\text{off}}(t)$. No transient signal can be observed for the single image acquisition since the noise level of $10\%$ overcomes the physical signal of $5\%$. However, the normalization $I_{\text{on}}(t)/I_{\text{off}}(t)$ for each delay $t$ increases the SNR by almost two orders of magnitude which reveals the physical transient signal due to the deflection of the electrons by the plasma generated after the photoexcitation.
When varying the flux of the electron, by varying the UV power, different transient dynamics are measured as shown in Fig. 4e). The transient signals are fitted with equation 4 to extract the temporal width of the electron beam $\sigma$. The temporal resolution gradually decreases as the electron flux increases.
The experimental results are compared to the mean-field calculation of the pulse length $l$ through the second-order differential equation Siwick et al. (2002):
$\frac{\text{d}^{2}l}{\text{d}t}=\frac{Ne^{2}}{m_{e}\epsilon_{0}\pi r^{2}}\left(1-\frac{l}{\sqrt{l^{2}+4r^{2}}}\right).$ (5)
With $N$, the number of electrons per bunch, $e$ and $m_{e}$ the electron charge and mass respectively, $\epsilon_{0}$ the vacuum permittivity, and $r$ the electron beam size.
The electron pulse duration is defined by $\Delta t=lv$, with the electron pulse velocity depending on the electron acceleration $U$ through :
$v=\sqrt{\frac{2eU}{m_{e}}}.$ (6)
To work with the optimized configuration, we study the simulated temporal resolution depending on parameters that can be adjusted in the lab. Firstly, Fig. 5a) shows the electron pulse duration along the bunch’s propagation distance for different electron fluxes from the numerical solution of Eq. 5. It suggests that the optimum distance to place the sample from the electron source is around $15$ cm to keep the temporal resolution below $1$ ps with a decent flux of a few thousand electrons per pulse.
Then, Fig. 5b) shows the small variation of the pulse duration as a function of the initial pulse duration, going from 1 fs to 1 ps. On the other hand, the pulse duration at 15 cm away from the electron source increases significantly when reducing the electron energy, mainly due to the longer time it takes for the electron to reach the sample’s position. We extract that a few tens of keV electron energy is sufficient to keep the temporal resolution below $1$ ps, while the initial pulse duration is not a sensitive parameter.
Lastly, Fig. 5c) shows the increase of temporal broadening with the decrease of the electron beam size. The broadening becomes significant for a beam size of around $500$ $\mu$m. As opposed to the Transmission Electron Microscope (TEM), the electron beam is focused on the electron detector to probe the sample’s reciprocal space. Thus, the electron beam size is relatively large at the sample position in the UED setup - estimated at $400$ $\mu$m. After the sample, the electron beam size shrinks down to $100$ nm without altering the temporal resolution. Fig. 5c) also displays the expected increase of the pulse duration as a function of the electron flux for a constant initial pulse duration of 100 fs. The time resolution is kept under 1 ps for flux lower than $5\cdot 10^{3}$ electron/pulse and increases only slightly for low flux (under 500 electron/pulse). The experimental values extracted from the fit of Fig. 4e) are represented with the blue star markers. While the trend is consistent with the simulation, the temporal resolution is shifted by $1$ ps. The simulation doesn’t include the extreme space charge after the photo-emission when the electron bunch is accelerated in the cathode. In this region, a large Coulomb repulsion is expected within the dense electron bunch. However, we only expect a higher initial pulse duration which doesn’t affect its dynamic after the electron source as shown in Fig. 5b).
Figure 5: a) Simulated pulse duration for different electron fluxes along the propagation distance for a $r=200\,\mu$m accelerated at $U=40$ keV with an initial pulse duration of 100 fs. b) Pulse duration at the sample’s position (15 cm) as a function of the initial pulse duration $t_{0}$ (top axis), and electron energy $U$ (bottom axis). c) Pulse duration at sample’s position as a function of electron flux $N$, and electron beam size $r$. The experimental values are marked with blue stars for the pulse duration as a function of the number of electrons per pulse from Fig. 4e). For b) and c) the simulations are done by keeping $t_{0}=100$ fs, $r=200\,\mu$m, $U=40$ keV, and $N=10^{\prime}000$ when not varying.
@@ IV Realistic situation
The photo-induced plasma generation shows already a striking difference between the conventional and shot-to-shot acquisition methods. However, the intensity distribution with diffraction patterns relies on the sample geometry. Photo-induced thermal and mechanical effects can induce instability or artifacts. In Sec. III.2 we already showed that these effects are stable over time. In this section, we confirm the capabilities of the shot-to-shot acquisition method on a time-resolved experiment on Graphite.
Ultrafast electron diffraction experiment on Graphite has been already extensively studied Carbone et al. (2008); Chatelain et al. (2012); Stern et al. (2018). To account for the sensitivity of our acquisition techniques, we use a low absorbed excitation energy of 0.8 mJ/cm2 on a 50 nm thick flake of natural graphite. The diffraction patterns are acquired over 24 hours, for an exposure of 22 min per delay (11 min for $I_{\text{on}}(t)$ and $I_{\text{off}}(t)$ respectively). To overcome the slight zone axis misalignment and gain SNR, we averaged the diffraction patterns along the 6-fold symmetry axis, as shown in Fig. 6c).
To observe the photo-induced effect we normalize the excited diffraction patterns by the one at rest and subtract this by the same ratio averaged along delays before time zero, $t_{0}$, when the pump pulse arrives after the probe one. The subtraction removes the constant thermal effect. It can be written as:
@@
The transient dynamics account for the sensitivity of the setup qualitatively. Fig. 6d) shows the intensity variation for the 100100{100}100 Bragg peak family and the diffuse scattering at the $K$-point and around the $\Gamma_{{110}}$ Bragg peak highlighted in the diffraction pattern Fig 6c). The low excitation energy induces a diffraction decrease of only $0.1\%$ due to the Debye-Waller effect. We can still resolve the transition between the fast initial drop from the generation of strongly coupled optical (SCOP) phonon to the slower one from their decay to acoustic phonon at 1 ps. This is consistent with the fast increase of the diffuse scattering intensity at the K-point due to the generation of $A_{1}^{\prime}$ optical phonon. Its population starts decreasing at 2 ps as it scatters with lower energy near $\Gamma$ momentum acoustic phonons as suggested by the transient dynamics of diffuse scattering around the $\Gamma_{110}$ Bragg peak family. These results are consistent with the previous work of Stern et al Stern et al. (2018) in which the photoexcitation energy is 15 times higher than the one used here.
Figure 6: a) and b) are the normalized diffraction patterns as equation 7 and 8 respectively, averaged over a time delay, $t$, between 1 and 2 ps. c) 6-fold averaged diffraction patterns of Graphite, corresponding to $I_{\text{on}}(t=-2.4\text{ ps},\bm{q})$. d) Electron count variation for specific regions of interest highlighted in panel c) corresponding to the Bragg peak family 100100{100}100 and the diffuse scattering at reduced momentum $K$ and around the $\Gamma$ point.
@@ V Conclusion
Similar to ultrafast optical spectroscopy, the shot-to-shot acquisition method improves significantly the signal-to-noise ratio for the study of transient dynamics in solids and molecules. By phase locking a high acquisition rate direct electron detector, the noise reaches the shot noise limit. By introducing a photoexcitation pulse, we observe thermal and mechanical fluctuations that are stable over time and don’t introduce additional white noise. The signal can be improved by increasing the electron flux with the only limitations being the temporal resolution and the detection saturation. The first can be overcome with temporal compression while the latter is lifted in the low scattering region.
The simulation of the temporal broadening indicates the right balance between electron flux and temporal resolution. The high sensitivity allows the investigation of ultrafast lattice dynamics after a low photo-excitation perturbation. It avoids thermal effects that can interfere with coherent effects or initiate phase transition.
Shot-to-shot acquisition schemes can be implemented in other electronic spectroscopy methods such as ultrafast Lorentz TEM, time-resolved electron energy loss spectroscopy (tr-EELS), time and angle-resolved photoemission spectroscopy (tr-ARPES), and others. It paves the way for the investigation of sensitive materials in the low perturbation regime.
@@ Appendix A Quantum detection efficiency
When an electron collides with a detector’s pixel, charges are displaced and a signal is acquired. If it reaches a given threshold, the detector counts one electron. The threshold is factory-calibrated for different electron energies.
By measuring the electron flux for different UV power, we can study the quantum detection efficiency defined as
$QDE=\frac{n_{\text{el}}}{n_{\text{ph}}},$ (9)
where $n_{\text{el}}$ is the number of acquired electrons per pulse directly measured by the direct electron detector for one second exposure time. $n_{\text{ph}}=P/fE_{\text{ph}}$ is the number of photons per pulse for a repetition rate $f=20$ kHz, a photon energy $E_{\text{ph}}=4.66$ eV and beam power $P$ measured with a power meter.
Fig. 7a) shows the acquired electron, $n_{\text{el}}$ and the quantum detection efficiency, $QDE$, as a function of the input UV power. The latter reveals a plateau from 1 $\mu$W and decreases $10^{3}$ $\mu$W when the flux reaches $10^{3}$ electrons/pulse.
The flux saturation can be due to the space charge and thermal effect of the photocathode. Another more probable explanation is the detection saturation. The detector counts one electron for a given pixel when its signal reaches a threshold. When the signal is higher than the threshold for longer than 200 ns, the detector counts another electron. This retrigger unit enables high electron flux measurement for continuous beam, but for a short pulsed beam of time duration $\ll$ 200 ns, the electron bunch is dense enough so that the detector cannot acquire more than one electron per pixel per pulse. Fig. 7b) illustrates this by showing the electron count for one pixel as a function of the input UV power. From one electron/pulse/pixel, the acquired count starts saturating. It is not favored to use a high pulsed electron flux with this kind of direct electron detector, as artifacts can arise from the detection saturation.
Figure 7: (a) Electron flux and quantum detection efficiency (Eq. 9) as a function of the input UV power for an electron energy of 30 keV. (b) Acquired electron per pulse for one selected pixel as a function of the UV power.

### v2

@@ III.2 Noise level with pump
In the presence of pump pulses, other noise sources, such as thermal and mechanical fluctuations of the thin film, should be considered. To verify how it affects real measurement, we focus on a pump-probe experiment in which the probe arrives before the pump. The reversible signal after photo-excitation won’t interfere with our analysis, but only the long-lived thermal and mechanical fluctuations.
We take a 50 nm thick sample of natural graphite as a studied specimen whose diffraction patterns are shown in Fig. 5c). We photoexcite the sample with an absorbed energy density of 0.8 mJ/cm2 at a wavelength around 800 nm, similar to experiments previously reported Carbone et al. (2008); Chatelain et al. (2014); Stern et al. (2018). However, in this section, we analyze the signal before the arrival time of the photoexcitation pulse, when the sample is at equilibrium, which should be identical to the diffraction patterns without photoexcitation.
We acquire a set of $I_{\text{on}}^{(i)}$ and $I_{\text{off}}^{(j)}$ images corresponding to a ten-second exposure of graphite’s diffraction patterns when it is photoexcited for a time window $i$, and when it is at rest for a time window $j$, respectively. For each diffraction image, we used a mask to set the unscattered electron beam to zero, such that we compare only the fluctuation of the scattered electron. Despite a protection coating on the detector, the pump beam scattered in the chamber creates abnormally high-intensity pixels only visible on the $I_{\text{on}}$ images. To remove them, we compare each image $I_{\text{on}}^{(i)}$ to $I_{\text{off}}^{(i)}$. The pixels in $I_{\text{on}}^{(i)}$ respecting the conditions $I_{\text{on}}^{(i)}/I_{\text{off}}^{(i)}>n$ and $I_{\text{on}}^{(i)}/I_{\text{off}}^{(i)}<1/n$ are set to their value in $I_{\text{off}}^{(i)}$. For a cutting parameter $n=2$, around $1.8\%$ of the pixels are neutralized.
Fig. 3a) shows the correlation between $I_{\text{on}}^{(i)}$ and $I_{\text{off}}^{(i)}$ taken shot-to-shot, and between two diffraction patterns $I_{\text{on}}^{(i)}$ and $I_{\text{on}}^{(j)}$ taken a few seconds apart, that is, $i\neq j$. The pulse-to-pulse images are strongly correlated compared to the two single images, especially at low electron count. We can characterize the correlation between the two images with the Pearson coefficient:
@@
where the sum is over all the pixels $x$, and $\langle I\rangle$ corresponds to the mean over the pixels. Fig. 3b) shows the Pearson coefficient for a set of integrated images taken at different times. The correlation between $I_{\text{on}}^{(i)}$ and $I_{\text{off}}^{(i)}$ images taken shot-to-shot stays close to 1, while the one between $I_{\text{on}}^{(0)}$ and $I_{\text{on}}^{(j)}$ decreases and fluctuates with the time between exposure shots. The average correlation between $\langle I_{\text{on}}(i),I_{\text{off}}(i)\rangle_{i}$ is $1-6.8\cdot 10^{-4}$, with a standard deviation of $1.6\cdot 10^{-5}$. It suggests that the shot-to-shot acquisition mode effectively removes thermal and mechanical fluctuations.
@@
Fig. 3c) shows the pixel distribution of the normalization $I_{\text{on}}/I_{\text{off}}$ as a function of exposure time. We observe that it converges towards a Gaussian distribution with decreasing variance, reaching $0.01$ with a mean of $0.99$. It confirms the robustness of the correlation between the diffraction pattern of the unpumped and pumped graphite.
@@ IV Realistic situation
The analysis of the noise origins showed that the SNR increased by one order of magnitude for the shot-to-shot acquisition with respect to the conventional technique. The correlation between the pumped and the unpumped data also showed that this acquisition method negates mechanical fluctuations in a pulse-to-pulse regime. We now focus on time traces after photoexcitation for two different practical cases: high-fluence pump on a copper grid generating a cloud of free electrons, which forms an electric field deviating the electron beam; and a well-studied sample, graphite, perturbed with a low fluence.
IV.1 Plasma generation
We use a 1000 mesh copper grid and photoexcite at a wavelength of 800 nm with a fluence of 2 mJ/cm2. Fig. 4a) and b) show the unfocused electron beam that went through the copper grid in the absence, $I_{\text{on}}(t)$, and the presence, $I_{\text{off}}(t)$, of the pump, respectively, at time $t$ after the photoexcitation. Thanks to the shot-to-shot acquisition detection, the normalization between Fig. 4a) and b), $I_{\text{on}}/I_{\text{off}}$, shown in Fig. 4d), clearly reveals the pattern of the plasma after the photoexcitation.
Fig. 4c) shows the total count in a region of interest defined as a $6\times 6$ pixels highlighted by the square in Fig. 4d). Each delay corresponds to an exposure time $20\cdot 10^{3}$ pulses for both $I_{\text{on}}(t)$, and $I_{\text{off}}(t)$. No transient signal can be observed for the single image acquisition since the noise level of $>10\%$ overcomes the physical signal of $5\%$. However, the normalization $I_{\text{on}}(t)/I_{\text{off}}(t)$ for each delay $t$ increases the SNR by almost two orders of magnitude, which reveals the physical transient signal due to the deflection of the electrons by the plasma generated after the photoexcitation.
Figure 4: Transient signal due to plasma generation after photoexcitation on a copper grid. Zoom in on the unfocused electron beam that went through the copper grid after the photoexcitation photon pulse in a) the presence of the pump $I_{\text{on}}$, b) the absence of the pump $I_{\text{off}}$, and d) the normalization between each other $I_{\text{on}}/I_{\text{off}}$. c) Comparison between the transient signals measured for a region of interest in the center of the electron beam $I_{\text{on}}(t)$, $I_{\text{off}}(t)$, and $I_{\text{on}}(t)/I_{\text{off}}(t)$.
IV.2 Graphite
The photo-induced plasma generation already shows a striking difference between the conventional and shot-to-shot acquisition methods. However, the intensity distribution with diffraction patterns relies on the sample geometry. Photo-induced thermal and mechanical effects can induce instability or artifacts. In Sec. III.2 we already showed that these effects are stable over time. In this section, we confirm the capabilities of the shot-to-shot acquisition method on a time-resolved experiment on graphite.
Ultrafast electron diffraction on graphite has already been extensively studied Carbone et al. (2008); Chatelain et al. (2012); Stern et al. (2018). To account for the sensitivity of our acquisition techniques, we use a low absorbed excitation energy of 0.8 mJ/cm2 on a 50 nm thick flake of natural graphite. The diffraction patterns are acquired over 24 hours, for an exposure of 22 min per delay (11 min for $I_{\text{on}}(t)$ and $I_{\text{off}}(t)$ respectively). To overcome the slight zone axis misalignment and gain SNR, we averaged the diffraction patterns along the 6-fold symmetry axis.
To observe the photo-induced effect, we normalize the excited diffraction patterns by the one at rest and subtract this by the same ratio averaged along delays before time zero, $t_{0}$. This subtraction removes the constant thermal effect. It can be written as:
@@ IV.2 Graphite
The transient dynamics account for the sensitivity of the setup qualitatively. Fig. 5d) shows the intensity variation for the 100100{100}100 Bragg peak family and the diffuse scattering at the $K$-point and around the $\Gamma_{{110}}$ Bragg peak highlighted in the diffraction pattern Fig 5c). The low excitation energy induces a diffraction decrease of only $0.1\%$ due to the Debye-Waller effect. We can still resolve the transition between the fast initial drop from the generation of strongly coupled optical (SCOP) phonon to the slower one from their decay to acoustic phonon at 1 ps. This is consistent with the rapid increase of the diffuse scattering intensity at the K-point due to the generation of $A_{1}^{\prime}$ optical phonon. Its population starts decreasing at 2 ps as it scatters with lower energy near $\Gamma$ momentum acoustic phonons as suggested by the transient dynamics of diffuse scattering around the $\Gamma_{110}$ Bragg peak family. These results are consistent with the previous work of Stern et al Stern et al. (2018) in which the photoexcitation energy is 15 times higher than the one used here.
Figure 5: a) and b) are the normalized diffraction patterns as equation 4 and 5 respectively, averaged over a time delay, $t$, between 1 and 2 ps. c) 6-fold averaged diffraction patterns of graphite, corresponding to $I_{\text{on}}(t=-2.4\text{ ps},\bm{q})$. d) Electron count variation for specific regions of interest highlighted in panel c) corresponding to the Bragg peak family 100100{100}100 and the diffuse scattering at reduced momentum $K$ and around the $\Gamma$ point.
@@ V Conclusion
Similar to ultrafast optical spectroscopy, the shot-to-shot acquisition method significantly improves the signal-to-noise ratio for the study of transient dynamics in solids and molecules. By phase locking a high acquisition rate direct electron detector, the noise reaches the shot noise limit. By introducing a photoexcitation pulse, we observe thermal and mechanical fluctuations that are stable over time and don’t introduce additional white noise.
The shot-to-shot acquisition technique revealed a marked improvement over conventional methods. The investigation of plasma generation and transient dynamics in graphite diffraction patterns demonstrates a time resolution of approximately 1 ps, achieved with an electron flux of just 1000 electrons per pulse. In the absence of a pulse compression scheme, the temporal resolution varies with the electron flux, improving at lower flux and degrading as the flux increases.
The high sensitivity of the setup enables the investigation of ultrafast lattice dynamics following low-level photoexcitation. In this regime, the lattice temperature increases by only a few tens of degrees, allowing the observation of dynamics driven by fluctuations in electronic, magnetic, or collective excitations.
This acquisition technique offers sensitivity comparable to, or exceeding, that of other UED configurations operating at high repetition rates Freelon et al. (2023) or using high electron flux Chatelain et al. (2012), while providing enhanced sensitivity in the low-perturbation regime. Integrating this approach with a higher repetition-rate laser would increase the signal by one to two orders of magnitude, without compromising temporal resolution or beam quality.
Shot-to-shot acquisition schemes can be implemented in other electronic spectroscopy methods, such as ultrafast transmission electron microscopy (UTEM)Zewail (2010), ultrafast Lorentz TEMTengdin et al. (2022), time-resolved electron energy loss spectroscopy (tr-EELS)Carbone, Kwon, and Zewail (2009), time- and angle-resolved photoemission spectroscopy (tr-ARPES)Na et al. (2019), and others. This paves the way for investigating sensitive materials in the low-perturbation regime.
Acknowledgements.
We acknowledge the useful discussion with the Dectris team, especially Jan Vávra, who helped us install the detector.
We acknowledge support from the SNF grant 200331. B.W. acknowledges support from the EU H2020 research and innovation program under the Marie Skłodowska-Curie Actions grant agreement 801459 (FP-RESOMUS). This project has also received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement no. 871124 Laserlab-Europe.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@ Appendix A Quantum detection efficiency
When an electron collides with a detector’s pixel, charges are excited, and a current is measured. If it reaches a given threshold, the detector counts one electron. The threshold is factory-calibrated for different electron energies.
We are interested in finding the optimal working conditions of the electron beam. One way to characterize the electron beam is the quantum efficiency, defined as
$QE=\frac{n_{\text{el}}}{n_{\text{ph}}},$ (6)
where $n_{\text{el}}$ is the number of acquired electrons per pulse directly measured by the direct electron detector for one-second exposure time. $n_{\text{ph}}=P/fE_{\text{ph}}$ is the number of incident photons per pulse for a repetition rate $f=20$ kHz, a photon energy $E_{\text{ph}}=4.66$ eV and beam power $P$ measured with a power meter.
Fig. 6a) shows the acquired electron, $n_{\text{el}}$ and the quantum efficiency, $QE$, as a function of the input UV power. We observe three stages with a non-linear increase of the $QE$ at $P<1\mu$W, a plateau between 1 $\mu$W and 1 mW, and a decrease of the $QE$ above 1 mW. The first stage corresponds to an extremely low dose rate, below one electron per pulse, so undetectable by the detector, while the last stage corresponds to a flux saturation. The latter can arise from a saturation in the photoemission due to the space charge and thermal effects of the photocathode, the Child-Langmuir limit, or from a saturation of the detection. The detector counts one electron for a given pixel when its signal reaches a threshold. When the signal is higher than the threshold for longer than 200 ns, the detector counts another electron. This retrigger functionality enables high electron flux measurement for continuous beam, but for a short pulsed beam of time duration $\ll$ 200 ns, the electron bunch is too dense, such that the detector cannot acquire more than one electron per pixel per pulse. Fig. 6b) illustrates this by showing the electron count for one pixel as a function of the input UV power. From one electron/pulse/pixel, the acquired count starts saturating. It is not favored to use a high pulsed electron flux with this kind of direct electron detector, as artifacts can arise from the detection’s saturation.
As shown in Fig. 6, the optimal UV power range for electron photoemission lies between 1 $\mu$W and 1 mW, where the measured quantum efficiency reaches a maximum of approximately $4\cdot 10^{-8}$ electrons per photon.
Figure 6: (a) Electron flux and quantum efficiency (Eq. 6) as a function of the input UV power for an electron energy of 30 keV. (b) Acquired electrons per pulse for one selected pixel as a function of the UV power.